# Streaming Private Continual Counting via Binning

### v1

@@ I-D Related work
More recently, and independently of our work, Dvijotham, McMahan, Pillutla, Steinke, and Thakurta [9], identified the same problem and designed a near-optimal solution with space complexity $\Theta(\log^{2}n)$ and the same maximum squared error up to an additive $o(1)$ term. But they only prove this near-optimality with respect to the maximum squared error, not the mean squared error. Their solution is based on approximating $L$ and $R$ by Toeplitz matrices $\hat{L}$ and $\hat{R}$ where the values on subdiagonals of $\hat{L}$ satisfy a recursion that facilitates computing $L\mathbf{z}$ in low space. Their technique could potentially be extended to factorizations of different $A$, but their existing analysis only holds for the lower-triangular all-1s matrix and relies on all matrices involved being Toeplitz. Our approach does not need any assumption of Toeplitz structure and is arguably simpler. It has a weaker theoretical space bound, but our empirical work shows that it works well with much less space, suggesting that our analysis is not tight.
@@ II-A1 Linear algebra
For a matrix $A\in\mathbb{R}^{m\times n}$, we let $A_{i,j}\in\mathbb{R}$ for $i\in[m],j\in[n]$ denote an entry. In the case where $[a,b]\subseteq[n]$ and $A_{i,j}$ assumes the same value for all $j\in[a,b]$, we will abuse notation and let $A_{i,[a,b]}$ represent this scalar. We reserve the notation $\lvert A\rvert$ for the matrix with entries $\lvert A\rvert_{i,j}=\lvert A_{i,j}\rvert$. $\|A\|_{2}$ denotes the operator norm of a matrix and if $A$ is invertible then $\kappa(A)=\|A\|_{2}\|A^{-1}\|_{2}$ defines the condition number of $A$. Whenever a matrix product $AB$ is given without specifying the dimensions of $A\in\mathbb{R}^{m\times n}$ and $B\in\mathbb{R}^{m^{\prime}\times n^{\prime}}$, we are implicitly imposing the restriction $n=m^{\prime}$.
@@ II-A3 Differential privacy
For $(\varepsilon,\delta)$-DP, it is known [5] that choosing $L=R=\sqrt{A}$ gives a continual counting mechanism with optimal mean squared error up to lower-order vanishing terms across all factorizations. The same factorization also improved the best known maximum (expected) squared error [6]. For this error, it was recently shown to be optimal up to the leading constant, ignoring lower-order terms, across all lower-triangular and Toeplitz factorizations [9]. We call the corresponding matrix $B=\sqrt{A}$ the Bennett matrix [17]. Given these strong error bounds, choosing the factorization $L=R=B$ would seem to be the obvious choice in practical settings.
@@ II-D Space matters
III Approximation by binnning
Given a factorization $LR=A$ where $L$ is lower-triangular and invertible, we are looking to approximate it, producing a new factorization $\hat{L}\hat{R}=A$ where $\operatorname{MeanSE}(\hat{L},\hat{R})\approx\operatorname{MeanSE}(L,R)$, $\operatorname{MaxSE}(\hat{L},\hat{R})\approx\operatorname{MaxSE}(L,R)$, and $\hat{L}$ has space complexity $o(n)$. We will approach the problem by expressing $\hat{L}$ as the sum of $L$, and a perturbation to each entry, $P$. The big picture idea is that, for some class of $L$, perturbing its elements a little might be sufficient to achieve sublinear space. Furthermore, if the perturbation $P$ is small enough and preserves the invertibility of $L$, then the requisite norms of $\hat{L}$ and $\hat{R}=\hat{L}^{-1}A$ will be close to those of $L$ and $R$. We attack this problem next.
@@ III-A Small perturbations preserve our norms of interest
We have that $\|L+P\|_{F}\leq\|L\|_{F}+\|P\|_{F}\leq\|L\|_{F}+\|\eta\lvert L\rvert+\mu J\|_{F}\leq(1+\eta)\|L\|_{F}+\mu\|J\|_{F}=(1+\eta)\|L\|_{F}+\mu n$ where $J\in\mathbb{R}^{n\times n}$ is the all-1s matrix, proving the statement. The first and last inequalities are cases of the triangle inequality, whereas the second step uses the definition of a $(\eta,\mu)$-perturbation and the fact that Frobenius norm only depends on the absolute value of each entry. The statement for $\|L+P\|_{2\to\infty}$ follows from an identical argument, except that $\|J\|_{2\to\infty}=\sqrt{n}$. ∎
@@
$\frac{\operatorname{MaxSE}(\hat{L},\hat{R})}{\operatorname{MaxSE}(L,R)}=\frac{\|\hat{L}\|_{2\to\infty}^{2}\|\hat{R}\|_{1\to 2}^{2}}{\|L\|_{2\to\infty}^{2}\|R\|_{1\to 2}^{2}}$
@@ III-B Binned matrices are space (and time) efficient
To prove Lemma 7 and 8, we begin by proving intermediate results for how the computation $L\mathbf{z}$ can be performed efficiently using an internal state of size $\lvert\mathcal{B}\rvert$.
@@ III-C Perturbing matrices into binned matrices
While binned matrices are nice, the issue is that most matrices do not admit very interesting binnings. The Bennett matrix for example only admits the trivial binning. Our intention is therefore to approximate a matrix $L$ by a matrix $\hat{L}$ that admit non-trivial binnings, but are still “close” to $L$.
@@
At this point, we have shown that $\mathcal{B}$-approximated matrices derived from running Algorithm 1 have space and time complexity linear in the binning size. We show next that, for a certain class of matrices, Algorithm 1 can produce binnings achieving a good trade-off between binning size and error when applied to the left factor in a factorization.
@@
Now consider the case where $\tau\geq\min_{i\geq j}L_{i,j}$ and line 17 evaluates to true for a row $i\in[n]$. In this case every interval except the left-most one of the form $[1,b]$ is guaranteed to satisfy the same condition. However, since $\hat{L}_{i,[1,b]}$ will be set to the average of two positive values less than or equal to $\tau$, we instead get that $\lvert P_{i,j}\rvert\leq\lvert\hat{L}_{i,j}-L_{i,j}\rvert\leq\tau$ for $j\in[1,b]$. Thus $P$ satisfies $\lvert P_{i,j}\rvert\leq(1/c^{2}-1)\lvert L_{i,j}\rvert+\tau$ for every $i,j\in[n]$.
Lastly, as invertibility of a lower-triangular matrix is decided by its diagonal, we also note that $\hat{L}$ is invertible, since the $\mathcal{B}$ produced by Algorithm 1 satisfies $[i,i]\in\mathcal{B}^{i}$, and so $\hat{L}_{i,i}=L_{i,i}$. ∎
@@ III-D Applications to square-root factorizations for counting
$=\sum_{i=0}^{k}\alpha^{k-i}\gamma(i)\gamma^{\prime}(k-i)\beta^{i}$
@@
$=\alpha^{k}\sum_{i=0}^{k}\bigg{(}\frac{\beta}{\alpha}\bigg{)}^{i}\gamma^{\prime}(i)\gamma^{\prime}(k-i)$
@@
Invoking Lemma 13 we get that $\|B_{\alpha,\beta}\|_{2}\leq\sum_{k=0}^{n-1}\lvert b_{k}\rvert\leq\frac{1}{1-\beta/\alpha}\sum_{k=0}^{n-1}\frac{\alpha^{k}}{\sqrt{k+1}}$ where the bound on $b_{k}$ comes from 2. If $\alpha=1$, then we have that $\frac{1}{\sqrt{k+1}}\leq\int_{k}^{k+1}\frac{\mathsf{d}x}{\sqrt{x}}$ for $k>0$ and thus $\|B_{\alpha,\beta}\|_{2}\leq\frac{1}{1-\beta}(1+\int_{1}^{n}\frac{\mathsf{d}x}{\sqrt{x}})=\frac{1}{1-\beta}(1+2\sqrt{n}-2)=\frac{2\sqrt{n}-1}{1-\beta}$. In the case where $\alpha<1$ we can write $\|B_{\alpha,\beta}\|_{2}\leq\frac{1}{1-\beta/\alpha}\sum_{k=0}^{n-1}\alpha^{j}\leq\frac{1}{(1-\beta/\alpha)(1-\alpha)}$. ∎
@@ IV-B Results for the Bennett matrix
This is not a computational fluke. Simply put, the optimality results in [5, 9] show that $B^{2}=A$ is asymptotically optimal, but on fixing $n$ one can hope to improve lower order terms for the error, and, as we show, even in sublinear space. We go on to show that this fact is demonstrably true up to at least $n=10^{4}$.
@@ V Conclusion and discussion
We have introduced a new framework for approximating factorization mechanisms based on a technique called “binning”. We have provided an algorithm that always produces a valid binning, and we show that for a class of factorizations $LR$, running it with suitable parameters will produce a new factorization $\hat{L}\hat{R}$ that runs in sublinear space at the cost of a small multiplicative increase in error. This class includes the asymptotically optimal, up to lower order terms, factorization of the counting matrix $L=R=\sqrt{A}$ in [6, 5, 17]. Empirically we show something stronger. For $n$ at least up to $10^{4}$, we can produce factorization with lower error in sublinear space. For weaker multiplicative guarantees, our empirical results suggest that our approach leads to logarithmic space complexity. Beyond this, we also show that our approach works once momentum and weight decay are introduced to the counting matrix.
There are many possible directions for future work. The main question we leave is the space efficiency possible by binning: our theoretical bounds on space and error do not appear to be tight in general, as demonstrated by our empirical work. An approach that relies on more global properties of the perturbation $P=\hat{L}-L$, rather than bounding each element, could be a way forward. Another question worth pursuing is what other linear query matrices (and their factorizations) can be worth studying under binning. This is in particular relevant to non-Toeplitz matrices, where the techniques in [9] are not applicable. It would also be of interest to investigate whether the binnings produced by Algorithm 1 can be substantially improved.
Finally, we mention that the bit complexity of space-efficient mechanisms that rely on real-valued noise could be worth a closer look. It seems straightforward to replace Gaussian noise by discrete Gaussian noise [22] in our mechanism: Since we only add random samples together there are no numerical issues and we can get high accuracy with only $O(\log n)$ bits per value. For other methods such as [9] it is less clear how many bits of precision are needed to maintain the privacy/utility trade-off.

### v2

@@ I Introduction
A version of this work has been accepted for publication in the IEEE Conference on Secure and Trustworthy Machine Learning (SaTML). The final version will be available on IEEE Xplore.
@@ I-D Related work
More recently, and independently of our work, Dvijotham, McMahan, Pillutla, Steinke, and Thakurta [9], identified the same problem and designed a near-optimal solution with space complexity $\Theta(\log^{2}n)$ and the same maximum squared error up to a $(1+o(1))$ multiplicative factor. They prove near-optimality with respect to the maximum squared error but do not state any results for mean squared error. Their solution is based on approximating $L$ and $R$ by Toeplitz matrices $\hat{L}$ and $\hat{R}$ where the values on subdiagonals of $\hat{L}$ satisfy a recursion that facilitates computing $L\mathbf{z}$ in low space. Their technique could potentially be extended to factorizations of different $A$, but their existing analysis only holds for the lower-triangular all-1s matrix and relies on all matrices involved being Toeplitz. Our approach does not need any assumption of Toeplitz structure and is arguably simpler. It has a weaker theoretical space bound, but our empirical work shows that it works well with much less space, suggesting that our analysis is not tight.
@@ II-A1 Linear algebra
For a matrix $A\in\mathbb{R}^{m\times n}$, let $A_{i,j}\in\mathbb{R}$ for $i\in[m],j\in[n]$ denote an entry. In the case where $[a,b]\subseteq[n]$ and $A_{i,j}$ assumes the same value for all $j\in[a,b]$, we will abuse notation and let $A_{i,[a,b]}$ represent this scalar. We reserve the notation $\lvert A\rvert$ for the matrix with entries $\lvert A\rvert_{i,j}=\lvert A_{i,j}\rvert$. $\|A\|_{2}$ denotes the operator norm of a matrix and if $A$ is invertible then $\kappa(A)=\|A\|_{2}\|A^{-1}\|_{2}$ defines the condition number of $A$. Whenever a matrix product $AB$ is given without specifying the dimensions of $A\in\mathbb{R}^{m\times n}$ and $B\in\mathbb{R}^{m^{\prime}\times n^{\prime}}$, we are implicitly imposing the restriction $n=m^{\prime}$.
@@ II-A3 Differential privacy
For $(\varepsilon,\delta)$-DP, it is known [5] that choosing $L=R=\sqrt{A}$ gives a continual counting mechanism with optimal mean squared error up to lower-order vanishing terms across all factorizations. The same factorization also improved the best known maximum expected squared error [6]. For this error, it was recently shown to be optimal across all lower-triangular Toeplitz factorizations [9]. We call the corresponding matrix $B=\sqrt{A}$ the Bennett matrix [17]. Given these strong error bounds, choosing the factorization $L=R=B$ would seem to be the obvious choice in practical settings.
@@ II-D Space matters
III Approximation by binning
Given a factorization $LR=A$ where $L$ is lower-triangular and invertible, we are looking to approximate it, producing a new factorization $\hat{L}\hat{R}=A$ where $\operatorname{MeanSE}(\hat{L},\hat{R})\approx\operatorname{MeanSE}(L,R)$, $\operatorname{MaxSE}(\hat{L},\hat{R})\approx\operatorname{MaxSE}(L,R)$, and $\hat{L}$ has space complexity $o(n)$. We will approach the problem by expressing $\hat{L}$ as the sum of $L$, and a perturbation to each entry $P$. The big picture idea is that, for some class of $L$, perturbing its elements a little might be sufficient to achieve sublinear space. Furthermore, if the perturbation $P$ is small enough and preserves the invertibility of $L$, then the requisite norms of $\hat{L}$ and $\hat{R}=\hat{L}^{-1}A$ will be close to those of $L$ and $R$. We address this problem next.
@@ III-A Small perturbations preserve our norms of interest
We have that
$\|L+P\|_{F}$ $\leq\|L\|_{F}+\|P\|_{F}$
$\leq\|L\|_{F}+\|\eta\lvert L\rvert+\mu J\|_{F}$
$\leq(1+\eta)\|L\|_{F}+\mu\|J\|_{F}$
$=(1+\eta)\|L\|_{F}+\mu n,$
where $J\in\mathbb{R}^{n\times n}$ is the all-1s matrix, proving the statement. The first and last inequalities are cases of the triangle inequality, whereas the second step uses the definition of a $(\eta,\mu)$-perturbation and the fact that Frobenius norm only depends on the absolute value of each entry. The statement for $\|L+P\|_{2\to\infty}$ follows from an identical argument, except that $\|J\|_{2\to\infty}=\sqrt{n}$. ∎
@@
$\operatorname{MaxSE}(\hat{L},\hat{R})/\operatorname{MaxSE}(L,R)$
$=\frac{\|\hat{L}\|_{2\to\infty}^{2}\|\hat{R}\|_{1\to 2}^{2}}{\|L\|_{2\to\infty}^{2}\|R\|_{1\to 2}^{2}}$
@@ III-B Binned matrices are space (and time) efficient
To prove Lemmas 7 and 8, we begin by proving intermediate results for how the computation $L\mathbf{z}$ can be performed efficiently using an internal state of size $\lvert\mathcal{B}\rvert$.
@@ III-C Perturbing matrices into binned matrices
While binned matrices are nice, the issue is that most matrices do not admit very interesting binnings. The Bennett matrix for example only admits the trivial binning. Our intention is therefore to approximate a matrix $L$ by a matrix $\hat{L}$ that has a non-trivial binning, but is still “close” to $L$.
@@
At this point, we have shown that $\mathcal{B}$-binned matrices derived from running Algorithm 1 have space and time complexity linear in the binning size. We show next that, for a certain class of matrices, Algorithm 1 can produce binnings achieving a good trade-off between binning size and error when applied to the left factor in a factorization.
@@
Now consider the case where $\tau\geq\min_{i\geq j}L_{i,j}$ and line 17 evaluates to true for a row $i\in[n]$. In this case every interval except the left-most one of the form $[1,b]$ is guaranteed to satisfy the same condition. However, since $\hat{L}_{i,[1,b]}$ will be set to the average of two positive values less than or equal to $\tau$, we instead get that $\lvert P_{i,j}\rvert=\lvert\hat{L}_{i,j}-L_{i,j}\rvert\leq\tau$ for $j\in[1,b]$. Thus $P$ satisfies $\lvert P_{i,j}\rvert\leq(1/c^{2}-1)\lvert L_{i,j}\rvert+\tau$ for every $i,j\in[n]$.
Lastly, as invertibility of a lower-triangular matrix is decided by its diagonal, we also note that $\hat{L}$ is invertible, since the $\mathcal{B}$ produced by Algorithm 1 satisfies $[i,i]\in\mathcal{B}^{i}$, and therefore $\hat{L}_{i,i}=L_{i,i}$. ∎
@@ III-D Applications to square-root factorizations for counting
$=\sum_{i=0}^{k}\alpha^{k-i}\gamma(i)\gamma^{\prime}(k-i)\beta^{i}\enspace.$
@@
$=\alpha^{k}\sum_{i=0}^{k}\bigg{(}\frac{\beta}{\alpha}\bigg{)}^{i}\gamma^{\prime}(i)\gamma^{\prime}(k-i)\enspace.$
@@
Invoking Lemma 13 we get that
$\|B_{\alpha,\beta}\|_{2}\leq\sum_{k=0}^{n-1}\lvert b_{k}\rvert\leq\frac{1}{1-\beta/\alpha}\sum_{k=0}^{n-1}\frac{\alpha^{k}}{\sqrt{k+1}}$
where the bound on $b_{k}$ comes from 2. For $\alpha=1$ we have
$\frac{1}{\sqrt{k+1}}\leq\int_{k}^{k+1}\frac{1}{\sqrt{x}}\mathsf{d}x$
for $k>0$ and thus
$\|B_{\alpha,\beta}\|_{2}$ $\leq\frac{1}{1-\beta}(1+\int_{1}^{n}\frac{1}{\sqrt{x}}\mathsf{d}x)$
$=\frac{1}{1-\beta}(1+2\sqrt{n}-2)=\frac{2\sqrt{n}-1}{1-\beta}\enspace.$
In the case where $\alpha<1$ we can write
$\|B_{\alpha,\beta}\|_{2}\leq\frac{1}{1-\beta/\alpha}\sum_{k=0}^{n-1}\alpha^{j}\leq\frac{1}{(1-\beta/\alpha)(1-\alpha)}\enspace.$
@@ IV-B Results for the Bennett matrix
This is not a computational fluke. Simply put, the optimality results in [5, 9] show that $B^{2}=A$ is asymptotically optimal, but on fixing $n$ one can hope to improve lower order terms for the error, and, as we show, even in sublinear space. In the particular case of the maximum squared error, $B^{2}=A$ is exactly optimal over Toeplitz factorizations [9], implying that our binned factorizations can achieve a lower error than any Toeplitz factorization of $A$. We go on to show that this fact is demonstrably true up to at least $n=10^{4}$.
@@ V Conclusion and discussion
We have introduced a new framework for approximating factorization mechanisms based on a technique we call “binning”. We give an algorithm that always produces a valid binning, and we show that, for a class of factorizations $LR$, calling it with suitable parameters produces a new factorization $\hat{L}\hat{R}$ running in sublinear space at the cost of a small multiplicative increase in error. This class includes the asymptotically optimal, up to lower order terms, factorization of the counting matrix $L=R=\sqrt{A}$ in [6, 5, 17]. Empirically we show something stronger. For $n$ at least up to $10^{4}$, we produce factorizations with lower error in sublinear space. For weaker multiplicative guarantees, our empirical results suggest that our approach leads to logarithmic space complexity. Beyond this, we also show that our approach works once momentum and weight decay are introduced to the counting matrix.
There are many possible directions for future work. The main question we leave is the space efficiency possible by binning: our theoretical bounds on space and error do not appear to be tight in general, as demonstrated by our empirical work. An approach relying on more global properties of the perturbation $P=\hat{L}-L$, rather than bounding each entry, could be a way forward. Another question worth pursuing is what other linear query matrices (and their factorizations) can be worth studying under binning. This is in particular relevant to non-Toeplitz matrices, where the techniques in [9] are not applicable. It would also be of interest to investigate whether the binnings produced by Algorithm 1 can be substantially improved.
Finally, we mention the bit complexity of space-efficient mechanisms that rely on real-valued noise. It seems straightforward to replace Gaussian noise by discrete Gaussian noise [22] in our mechanism: Since we only add random samples together there are no numerical issues and we can get high accuracy with only $O(\log n)$ bits per value, assuming that the exact prefix sums can be represented in $O(\log n)$ bits. For other methods such as [9] it is less clear how many bits of precision are needed to maintain the privacy/utility trade-off.